The speed of light in some other universe is 30 m/s. You want to cover a distance of 100 m in 10 s (as measured on your watch) at a constant speed.
a. How fast should you run?
b. How long did it take you to cover that distance according to an observer at rest with respect to the track? (Carry four sig figs throughout the calculation).
Answer #1

a)

v = d/ t = 100 / 10 = 10 m/s

======

b)

considering time dilation

t = to / sqrt ( 1 - v^2/c^2)

10 = to / sqrt ( 1 - 10^2/30^2)

to = 8.16497 s

=======